# Multi-Agent Path Finding under Limited Communication Range Constraint via Dynamic Leading

### v1

@@ Methods
3:     if $u.t\leq\zeta^{s}_{i}.\text{maxtime}$ then ${{\mathrm{lead}}}\leftarrow{{\texttt{false}}}$
@@
8:  return $({{\mathrm{lead}}}\ \|\ {{\mathrm{comm}}})$
SAPF MODULE
Single-Agent Path Finding (SAPF) finds a shortest reach-goal path for an agent. This planner searches actions on $\mathcal{G}$ that satisfy the ACOMM constraint and are collision-free. With enough runtime, the planner can explore all possible paths and so is probabilistically complete. If it fails, it returns the closest-to-goal path. Agents are assumed to have the same velocity ${{\mathrm{v}}}_{c}$. The inputs are the agent’s ID $\rho$, goal position $g_{\rho}$, a graph of the world sub-divisions $\mathcal{G}$, planned paths of previous agents $\zeta^{s}$, and the shortest path to the goal heuristic $\mathcal{H}$.
To initialize, a new graph $\mathcal{G}^{s}$ is created by adding to $\mathcal{G}$: (i) the current agent position $\zeta^{s}_{\rho}.{{\mathrm{end}}}$ and (ii) the goal $g_{\rho}$, each of which connect to the centroid of the sub-division to which they belong (Alg. 2:1). The variables $v_{g}$ and $v_{s}$ are the goal and current nodes (Alg. 2:2). $v_{s}$ is then added into ${{\mathrm{openList}}}$—a priority queue data structure, which sorts its elements by their cost-to-goal. closedSet and the closest-to-goal node $v_{\text{best}}$ are also initialized (Alg. 2:3).
The search loop (Alg. 2:4–18) is: a node with lowest cost is popped out from ${{\mathrm{openList}}}$ (Alg. 2:5); then we iterate through all its neighbors (Alg. 2:7–18) to find valid nodes, calculate or update their costs, parents, and timestep, then add them into openList.
@@
For a node $u$ to be valid, the move action from its parent $v$ must be collision-free and satisfy the ACOMM constraint. Both of these requirements are checked by the function ${\scriptstyle{{\textsc{ISNODEVALID}}}}()$ (Alg. 2:11,16).
@@
11:        $v_{n}.{{\mathrm{state}}}.{\scriptstyle{{\textsc{Insert}}}}({{\mathrm{pos}}})$; $d={\scriptstyle{{\textsc{Dist}}}}(v_{p}.{{\mathrm{state}}}_{j},v_{n}.{{\mathrm{state}}}_{j})$
@@
13:     $v_{n}.f\leftarrow v_{n}.g$ + $v_{n}.h$
14:     $\mathcal{T}_{{\mathrm{ma}}}.{\scriptstyle{{\textsc{Append}}}}(v_{n})$; $v_{p}\leftarrow v_{n}$
IsNodeValid() Function
This function (Alg. 3) checks the validity of action $\textit{Move}(v,u)$ for an agent $\rho$ starting at timestep $v.t$. For agent-collision detection, ${\scriptstyle{{\textsc{IsCollision}}}}()$ (Alg. 3:6) checks if any segmented point along the line $l_{vu}$ (connecting $v$ and $u$) collides with the planned paths in $\zeta^{s}$ (Alg. 3:5) at corresponding timesteps. If a collision occurs, the node is invalid.
IsCommAtGoal() Function
EXPAND MULTI-AGENT TREE
Function ${\scriptstyle{{\textsc{ExpandMATree}}}}()$ (Alg. 4) integrates the paths from the single agents $\zeta^{s}$ and adds the combined paths to expand the multi-agent tree $\mathcal{T}_{{\mathrm{ma}}}$. Each node on $\mathcal{T}_{ma}$ has an associated timestep and the agent positions are interpolated to make them match for each single agent path.
@@
The for loop (Alg. 4:2–3) collects all timesteps of the waypoints on the single paths and inserts into a priority queue ${{\mathrm{timeList}}}$. The second for loop (Alg. 4:4–14) goes through all elements in ${{\mathrm{timeList}}}$ to create new nodes and add valid nodes into the tree $\mathcal{T}_{ma}$. Each new node $v_{n}$ inherits the travel cost from its parent with timestep $t$ (Alg. 4:5). We then go to each agent’s path, interpolate to recover its position at time $t$ (Alg. 4:7–12), and insert it into the state of node $v_{n}$. If $t$ is larger than the max-timestep of the path $\zeta^{s}_{j}$ (Alg. 4:7) and the path reaches the goal, the final position of $\zeta^{s}_{j}$ is returned (Alg. 4:8). If the position is not the goal, the tree’s expansion stops (Alg. 4:9).
@@
If $t$ is smaller or equal to the last timestep $\zeta^{s}_{j}$, the agent’s position at $t$ is interpolated by function ${\scriptstyle{{\textsc{GetPosAtTime}}}}()$ (Alg. 4:10). The lines in Alg. 4:11–12 add the agent’s positions into the node state; we then update the travel cost and the heuristic. Finally, the new node $v_{n}$ is added into the MATree $\mathcal{T}_{{\mathrm{ma}}}$, and $v_{p}$ is reassigned to continue the expansion (Alg. 4:14).
Experiments and Results
@@
Experiments are conducted on five obstacle-rich environments (Fig. 4). Our planner’s performance is measured by three metrics: (1) success-rate, (2) runtime, and (3) per-agent travel distance. Our MA-DL approach is compared with two baselines: a centralized approach with composite states and a platooning leader-follower approach. All agents have 8 actions: to move one unit in the four cardinal directions and $\sqrt{2}$ unit in their diagonal directions. We also evaluate the planner’s robustness versus of runtime, goal configuration, and difficulty of environments.
@@
Experimental Setup
Baselines
We have implemented two baselines: Centralized approach with composite state (COMP) and Platooning Leader-Follower approach (PLF). COMP is selected because existing MAPF planners—which typically advance the motion of each agent in isolation of one another and then attempt to resolve conflicts as necessary—are not well suited to handle the communication constraint, as the challenge of ensuring that all agents are in constant communication would require an undue number of repair operations, causing such planners to struggle (Section 2). Meanwhile, PLF is chosen because it is a state-of-the-art approach to solve MALCR problem.
  • COMP grows a multi-agent tree via A* using informed by a heuristic of sum of shortest-to-goal paths from each agent. In joint-actions of single and diagonal moves, the diagonal ones are trimmed down to equal length with the single moves.
  • PLF also builds a multi-agent tree and using priority planning to grows the tree. At the root of tree expansion, the planning order is shuffled and so a random leader is chosen; planning order is then fixed for downstream nodes. The followers plan to follow another agent if its preceding agent reached its goal. If the current expansion fails after some iterations, the planning starts again at the root with a new random leader and corresponding planning order.
Environments and Instances
For each number of agents, one instance is generated on each environment map. So, there are total of 12000 instances for the experiments. The start and goal configurations are chosen randomly on alternate sides of the maps (except the ring environment), distributed with a rectangular area. The configurations are satisfied the TCOMM constraint. Communication between two agents are established if the distance between them is within 15 m of one another.
Env. Type 1: Random Forest
(Fig. 4) Obstacles with random shapes and sizes are distributed randomly occupying 10% of the environment’s area.
Env. Type 2: Office
(Fig. 4) The environments consists multiple rooms and hallways. The room has a fixed width and varying length from 9–13 m. There are three long hallways along the building, 2–3 short hallways connecting them. The hallway width varies from 7–9 m.
Env. Type 3: Waves
(Fig. 4) The environments have wave-like obstacles which are separated at regular distances. Gaps are placed in each wave with random widths. The number of waves is set to $10$.
Env. Type 4: Rings
(Fig. 4) The environments are featured by concentric rings with six breaks of random widths within 6–8 m. The separation between the rings is set to 8 m. For each instance, the starts are randomly placed at the center, and the goals are on one of four corners of the maps.
Env. Type 5: Maze
(Fig. 4) The environments consist of mazes generated by Kruskal’s algorithm with size of $14\times 14$. To ease in generation of the starting and goal configurations, boundary walls connecting to the top-most and bottom-most rows are removed.
Measuring Performance
We evaluate our framework against COMP and PLF baselines on 100 instances for each environment type and number of agent, measuring success-rate, average runtime and average distance traveled per agent. Planning is considered successful if all agents reach their goals within 5 s of runtime. Runs that exceed that planning ceiling are assigned a travel distance of 300 m. The framework is evaluated with variations over number of agents, goal configurations, runtime, and environment difficulty levels.
Computing Resources
The experiments ran on HOPPER, a computing cluster provided by GMU’s Office of Research Computing. Each planning instance is run single threaded, yet experiment were run in parallel across 48 cores on a 2.40GHz processor. Our code was developed in C++ and compiled with g++-9.3.0.
Results
Results when varying the agent number
@@ Results
Both baseline planners struggle as the number of agents is increased. PLF manages up to 5 agents in Rings Env. and only 3-4 agents in other environments. The heuristic of shortest-to-goal path sum guides COMP well to handle up 4 agents in Random Forest, Rings, and Waves Env. However, with greater than 6 agents, the high dimensionality of the search space means that planning cannot reach the goal in the allotted time and success rate quickly declines.
@@
Results when varying runtime
In this experiment, we run the planners with various runtime from 1 s to 512 s. We select the Maze Env., the most challenging Env. type for the experiments. The results are shown in Fig. 5.a, showing that success-rate reaches 100% as the runtime grows.
@@
Results when varying environment difficulty
To demonstrate the robustness of our framework, we run MA-DL on variations of the Rings environment type with three difficulty levels: easy, medium, and hard. The difficulty is controlled by the number of concentric rings, distance between consecutive rings, and number of breaks/gates on the rings. The features controlling difficulty are shown in Table I. An easy and hard environment are shown in Fig. 5.b.
Diff. Level Ring Count Ring-Distance Break Count
@@
TABLE I: Environment Difficulty Features
@@
Results with Different Goal Configurations
@@
Figure 6: Long-Thin vs Rectangle of Goal Configuration.
Completeness Analysis
@@
Ma et al. [11] proved the following theorem for the incomplete prioritized planning:
Theorem 1: Prioritized (fixed) planning with an arbitrary priority ordering is incomplete for MAPF in general.
Based on this theorem, fixed and platooning leader-follower approaches are incomplete planners. A fix leader-follower planner never changes the planning order, while a platooning planner only does so when members leave the team. However, while our MA-DL planner overcomes this limitation, it is also an incomplete algorithm. This is a consequence of the greedy nature of the single agent planner (SAPF), which always immediately returns upon finding a valid single agent path. We illustrate a scenario in Fig. 7 that MA-DL fails to solve: though the solution requires that agent $a_{3}$ select a longer path to allow other agents to pass by, the SAPF planner returns only the shortest path and so the team becomes stuck.
Future work could explore extending our MA-DL approach to achieve completeness. For this to be possible, the SAPF single-agent planner would need to be extended to eventually generate all possible paths when expanding each agent, a modification that would require careful consideration and detailed study so that it would not dramatically slow planning performance. In this research, our focus is on developing a fast and practical planner that can efficiently handle the MALCR problem, outperforming existing approaches. We plan to develop a complete version of the MA-DL planner in the future research.
@@
Conclusion
This paper introduces the MA-DL framework to handle the MALCR problem. The core advance of our approach is dynamic leading, which enables the dynamic reselection of the leading agent during path planning whenever progress stalls, overcoming a key limitation of state-of-the-art platooning approaches. We have tested our MA-DL planner in multiple environments with features such as obstacles-rich, narrow passages, non-convex spaces, and long hallways. The results demonstrate the robustness of MA-DL planner, capable of planning up to 25 agents within 5 seconds of runtime. In future work, we aim to develop a complete version of this planner. We also intend to design a new heuristic to improve the expansion of single agent trees. Furthermore, we plan to extend the framework to support robot teams with rich kinodynamic constraints, bringing it a step closer to the real-world applications.
ACKNOWLEDGMENT

### v2

@@ 4.2 Single-Agent Path Finding: The SAPF Module
3:     if $u.t\leq\zeta^{s}_{i}.\text{maxtime}$ then ${{\mathrm{lead}}}\leftarrow{{\texttt{false}}}$ //is this agent leading?
@@
8:  return $({{\mathrm{lead}}}\ \|\ {{\mathrm{comm}}})$ //leader can go without comms. constraint
The search loop (Alg. 2:3–17) is: a node with lowest cost is popped out from ${{\mathrm{openList}}}$ (Alg. 2:4); then we iterate through all its neighbors (Alg. 2:6–17) to find valid nodes, calculate or update their costs, parents, and timestep, then add them into openList.
For a node $u$ to be valid, the move action from its parent $v$ must be collision-free and satisfy the ACOMM constraint. Both of these requirements are checked by the function ${\scriptstyle{{\textsc{ISNODEVALID}}}}()$ (Alg. 2:10,15).
IsNodeValid() Function
This function (Alg. 3) checks the validity of action $\textit{Move}(v,u)$ for an agent $\rho$ starting at time $v.t$. For agent-collision detection, ${\scriptstyle{{\textsc{IsCollision}}}}()$ (Alg. 3:6) checks if any segmented point along the line $l_{vu}$ connecting $v$ and $u$ collides with the planned paths in $\zeta^{s}$ (Alg. 3:5) at corresponding timesteps. A node is invalid if a collision occurs.
@@
IsCommAtGoal() Function
4.3 The Expand Multi-Agent Tree Function: ExpandMATree
Function ${\scriptstyle{{\textsc{ExpandMATree}}}}()$ (Alg. 4) integrates the paths from the single agents $\zeta^{s}$ and adds the combined paths to expand the multi-agent tree $\mathcal{T}_{{\mathrm{ma}}}$. Each node on $\mathcal{T}_{ma}$ has an associated timestep and positions are interpolated to make them match for each single agent path.
@@ 4.3 The Expand Multi-Agent Tree Function: ExpandMATree
11:        $v_{n}.{{\mathrm{state}}}.{\scriptstyle{{\textsc{Insert}}}}({{\mathrm{pos}}})$; $d={\scriptstyle{{\textsc{Dist}}}}(v_{p}.{{\mathrm{state}}}_{j},v_{n}.{{\mathrm{state}}}_{j})$;
@@
13:     $v_{n}.f\leftarrow v_{n}.g$ + $v_{n}.h$; $\mathcal{T}_{{\mathrm{ma}}}.{\scriptstyle{{\textsc{Append}}}}(v_{n})$; $v_{p}\leftarrow v_{n}$
@@
The for loop (Alg. 4:2–3) collects all timesteps of the waypoints on the single paths and inserts into a priority queue ${{\mathrm{timeList}}}$. The second for loop (Alg. 4:4–13) goes through all elements in ${{\mathrm{timeList}}}$ to create new nodes and adds valid nodes into the tree $\mathcal{T}_{ma}$. Each new node $v_{n}$ inherits the travel cost from its parent with timestep $t$ (Alg. 4:5). We then go to each agent’s path, interpolate to recover its position at time $t$ (Alg. 4:7–12), and insert it into the state of node $v_{n}$. If $t$ is larger than the max-timestep of the path $\zeta^{s}_{j}$ (Alg. 4:7) and the path reaches the goal, the final position of $\zeta^{s}_{j}$ is returned (Alg. 4:8). If the position is not the goal, the tree’s expansion stops (Alg. 4:9).
@@
If $t$ is smaller or equal to the last timestep $\zeta^{s}_{j}$, the agent’s position at $t$ is interpolated by function ${\scriptstyle{{\textsc{GetPosAtTime}}}}()$ (Alg. 4:10). The lines in Alg. 4:11–12 add the agent’s positions into the node state; we then update the travel cost and the heuristic. Finally, the new node $v_{n}$ is added into the MATree $\mathcal{T}_{{\mathrm{ma}}}$, and $v_{p}$ is reassigned to continue the expansion (Alg. 4:13).
5 Experiments and Results
@@
Experiments are conducted on five obstacle-rich environments (Fig. 4). Our planner’s performance is measured by three metrics: (1) success-rate, (2) runtime, and (3) per-agent travel distance. Our MA-DL approach is compared with two baselines: a centralized approach with composite states and a platooning leader-follower approach. All agents have 8 actions: to move one unit in the four cardinal directions and $\sqrt{2}$ unit in their diagonal directions. We also evaluate the planner’s robustness versus runtime, goal configuration, and difficulty of environments.
5.1 Experimental Setup
5.1.1 Baselines
We have implemented two baselines: Centralized planning with a composite state (COMP) and Platooning Leader-Follower approach (PLF). COMP is selected because existing MAPF planners—which typically advance the motion of each agent in isolation of one another and then attempt to resolve conflicts as necessary—are not well suited to handle the communication constraint, as the challenge of ensuring that all agents are in constant communication would require an undue number of repair operations, causing such planners to struggle (Section 2). PLF (Platooning Leader-Follower) is chosen because it is a competitive, state-of-the-art baseline for MALCR.
  • COMP (centralized planning with a composite state) grows a multi-agent tree via $A^{*}$ with a heuristic that sums shortest-to-goal paths from each agent.
  • PLF also builds a multi-agent tree and uses priority planning to grow the tree. At the root of tree expansion, the planning order is shuffled and so a random leader is chosen; planning order is then fixed for downstream nodes. The followers plan to follow another agent if its preceding agent reached its goal. If the current expansion fails after some iterations, the planning starts again at the root with a new random leader and a corresponding planning order.
5.1.2 Environments and Instances
For each number of agents, one instance is generated on each environment map. So, there are total of 12000 instances for the experiments. The start and goal configurations are chosen randomly, subject to the TCOMM constraint, on alternate sides of the maps (except the ring environment), distributed with a rectangular area. Communication between two agents is established if the distance between them is within 15 m of one another.
Env. Type 1: Random Forest
Obstacles with random shapes and sizes are distributed randomly occupying 10% of the environment’s area.
Env. Type 2: Office
The environments consist multiple rooms and hallways. Each room has a fixed width and varying length from 9–13 m. There are three long hallways along the building, 2–3 short hallways connecting them. Hallway width varies from 7–9 m.
Env. Type 3: Waves
The environments have wave-like obstacles which are separated at regular distances. Gaps are placed in each wave with random widths. The number of waves is set to $10$.
Env. Type 4: Rings
The environments are featured by concentric rings with six breaks of random widths within 6–8 m. The separation between the rings is set to 8 m. For each instance, the starts are randomly placed at the center, and the goals are on one of four corners of the maps.
Env. Type 5: Maze
The environments consist of mazes generated by Kruskal’s algorithm with size of $14\times 14$. To ease in generation of the starting and goal configurations, boundary walls connecting to the top-most and bottom-most rows are removed.
5.1.3 Measuring Performance
We evaluate our framework against COMP and PLF baselines on 100 instances for each environment type and number of agents, measuring success-rate, average runtime and average distance traveled per agent. Planning is considered successful if all agents reach their goals within 5 s of runtime. Runs that exceed that planning ceiling are assigned a travel distance of 300 m. The framework is evaluated with variations over number of agents, goal configurations, runtime, and environment difficulty levels.
5.1.4 Computing Resources
The experiments ran on HOPPER, a computing cluster provided by GMU’s Office of Research Computing. Each planning instance was run single threaded, yet experiments were run in parallel across 48 cores on a 2.40GHz processor. Our code was developed in C++ and compiled with g++-9.3.0.
@@ 5.1.4 Computing Resources
5.2 Results
5.2.1 Results when varying the agent number
@@
Both baseline planners struggle as the number of agents increases. PLF manages up to 5 agents in Rings Env. and only 3-4 agents in other environments. The heuristic of shortest-to-goal path sum guides COMP well to handle up 4 agents in Random Forest, Rings, and Waves Env. However, with greater than 6 agents, the high dimensionality of the search space means that planning cannot reach the goal in the allotted time and success rate quickly declines.
@@
5.2.2 Results when varying runtime
In this experiment, we ran the planners with various runtime from 1 s to 512 s. We selected the Maze Env., the most challenging Env. type for the experiments. The results are illustrated in Fig. 5.a, showing that success-rate reaches 100% as the runtime grows.
@@ 5.2.2 Results when varying runtime
5.2.3 Results when varying environment difficulty
To demonstrate the robustness of our framework, we run MA-DL on variations of the Rings environment type with three difficulty levels: easy, medium, and hard. The difficulty is controlled by the number of concentric rings, distance between consecutive rings, and number of breaks/gates on the rings. The features controlling difficulty are shown in Table 1. An easy and a hard environment are illustrated in Fig. 5.b.
Difficulty Ring Ring Break
Level Count Distance Count
@@ 5.2.3 Results when varying environment difficulty
Table 1: Environment Difficulty Features
@@
Figure 6: Long-Thin vs Rectangle of Goal Configuration.
5.2.4 Results with Different Goal Configurations
@@
6 Completeness Analysis
@@
Ma et al. (?) proved the following theorem for the incomplete prioritized planning: Prioritized (fixed) planning with an arbitrary priority ordering is incomplete for MAPF in general. Based on this theorem, fixed and platooning leader-follower approaches are incomplete planners. A fixed leader-follower planner never changes the planning order, while a platooning planner only does so when members leave the team. However, while our MA-DL planner overcomes this limitation, it is also an incomplete algorithm. This is a consequence of the greedy nature of the single agent planner (SAPF), which always immediately returns upon finding a valid single agent path. We illustrate a scenario in Fig. 7 that MA-DL fails to solve: though the solution requires that agent $a_{3}$ selects a longer path to allow other agents to pass by, the SAPF planner returns only the shortest path and so the team becomes stuck.
@@
Future work could explore extending our MA-DL approach to achieve completeness. For this to be possible, the SAPF single-agent planner would need to be extended to eventually generate all possible paths when expanding each agent, a modification that would require careful consideration and detailed study so that it would not dramatically slow planning performance. In this research, our focus is on developing a fast and practical planner that can efficiently handle the MALCR problem, outperforming existing approaches. We plan to develop a complete version of the MA-DL planner in future research.
7 Conclusion
This paper introduces the MA-DL planner to handle the MALCR problem. The core advance of our approach is dynamic leading, which enables the dynamic reselection of the leading agent during path planning whenever progress stalls, overcoming a key limitation of state-of-the-art platooning approaches. We have tested our MA-DL planner in multiple environments with features such as obstacles-rich, narrow passages, non-convex spaces, and long hallways. The results demonstrate the robustness of MA-DL planner, capable of planning up to 25 agents within 5 seconds of runtime. In future work, we aim to develop a complete version of this planner. We also intend to design a new heuristic to improve the expansion of single agent trees. Furthermore, we plan to extend the framework to support robot teams with rich kinodynamic constraints, bringing it a step closer to real-world applications.
Acknowledgments